# Impact of data usage for forecasting on performance of model predictive control in buildings with smart energy storage

### v1

@@ 2.2 Forecasting task & performance evaluation
For the MPC scheme used, Eq. 2.1, forecasts of the following operational condition variables over the planning horizon $T$ are required at each time instance, $t$:
@@
This study investigates the impacts of data usage on the forecasting of these 4 types of operational variables. The accuracies of these forecasts are quantified using two error metrics, the normalised Mean Absolute Error (nMAE), and the normalised Root Mean Squared Error (nRMSE), given by Eqns 2 & 3,
$\frac{{\sum_{t=0}^{N-1}}\,\frac{1}{T}\sum_{\tau=1}^{T}\mathinner{\!\left\lvert f^{v}_{t,\tau}-v_{t+\tau}\right\rvert}}{\sum_{t=0}^{N-1}v_{t}}$ (2)
$\frac{{\sum_{t=0}^{N-1}}\sqrt{\frac{1}{T}\sum_{\tau=1}^{T}\left(f^{v}_{t,\tau}-v_{t+\tau}\right)^{2}}}{\sum_{t=0}^{N-1}v_{t}}$ (3)
@@
The operational performance achieved by the MPC scheme using a given set of prediction models is quantified by evaluating the objective specified in Eq. 2.1 with the simulated behaviour of the building energy system resulting from the use of the controller.
@@ 2.3 Cambridge Estates electricity use dataset
A dataset of historic building electricity usage measurements from a set of buildings across the Cambridge University Estates covering the period 2010 to 2019 langtry2023CambridgeUniversityEstates is used for the example multi-building energy system. The dataset consists of 10 years of electricity usage data from 30 buildings of varying use types in the University Estate, such as lecture blocks, offices, laboratories, and museums, alongside weather observations and grid electricity price and carbon intensity data. In all, the dataset contains the following variables: building electrical load data, weather data for Cambridge from the Met Office MIDAS dataset metoffice2022MIDASOpenUK (temperature and relative humidity) and renewables.ninja reanalysis model pfenninger2016LongtermPatternsEuropean ; staffell2016UsingBiascorrectedReanalysis (direct and diffuse solar irradiance), dynamic electricity pricing tariff data from Energy Stats UK energystatsuk2023HistoricalPricingData , grid electricity carbon intensity from the National Grid ESO Data Portal nationalgrideso2020HistoricGenerationMix , and temporal information including hour, day, and month indices, as well as daylight savings status. All data is available at hourly resolution. Further detail on the data sourcing and processing is available in reference langtry2023CambridgeUniversityEstates .
@@ 2.4.1 Simple neural models
Recent literature, zeng2023AreTransformersEffective , has shown that simple neural models using Direct Multi-Step forecasting (DMS), where all predictions over the forecast window are generated concurrently, can outperform complex, transformer-based models using traditional, Iterated Multi-Step forecasting (IMS), in which a single-step forecaster is applied iteratively to generate a multi-step forecast. For all three simple neural architectures investigated, DMS forecasting is used.
@@ 2.4.2 State-of-the-art machine learning models
The Temporal Fusion Transformer (TFT) model lim2021TemporalFusionTransformers , developed by Google, is an attention-based architecture that enables the fusion of data from multiple input sources to inform predictions. The neural structure contains features which allow for the learning of multiple underlying relationships across temporal scales, and the attention mechanism allows for interpretation of the model predictions, i.e. which data the model is exploiting to produce its forecasts. The model uses categorical covariates of date-time information, as well as temperature information for predicting building loads.
@@
Neural Hierarchical Interpolation for Time Series Forecasting (NHiTS) challu2023NHITSNeuralHierarchical is an MLP model which learns a set of basis functions at different frequencies that describe the underlying patterns in the training data, and produces forecasts by using hierarchical interpolation to combine predictions from the basis functions in a computationally efficient manner. It uses categorical covariates of date-time information.
@@
DeepAR is a Recurrent Neural Network (RNN) based model developed by Amazon salinas2020DeepARProbabilisticForecasting , which has been widely applied in a range of research areas. It is a probabilistic forecasting model, but for this study only the mean prediction is used. The model uses categorical covariates of date-time information.
@@ 3.1 Baseline prediction accuracy comparison
The prediction accuracy of each forecasting model trained using the full 8 years of available training data was evaluated first to provide a baseline for the model architectures in a setting without data limitations. For brevity, forecast accuracy results are discussed for the nRMSE metric only, however equivalent results were found with the nMAE metric. Figures 2a & 2b show that the simple Linear model achieves similar or better prediction accuracy (lower nRMSE values) compared to the complex, state-of-the-art models across all prediction variables. The Conv and ResMLP models provide similar accuracy when predicting building electrical loads, however both have significantly worse accuracy for the electricity price and carbon intensity prediction variables. Complex models achieve slightly better accuracy for electricity price and solar generation predictions, at most 8.5% and 4.2% lower nRMSE than the Linear model, achieved by NHiTS and TFT respectively. However, for some buildings the complex models exhibit very poor forecast accuracy for load predictions. These instances of poor accuracy are found to correlate with a measure of the similarity between the train and test datasets for building load, called the ‘Wasserstein similarity metric’, which is described in D and is used for the study of model reuse in the following section. Fig. 2c plots the relationship between prediction accuracy and data similarity for each model, and shows that complex models provide poor prediction accuracy when the training and test data are significantly different. Hence, simple neural models provide better prediction generalisation under changes in building load dynamics between the training and test data, which is highly advantageous for application to practical systems, as occupant driven load dynamics may change after system installation, e.g. due to a change of building use.
Simple neural models are able to provide analogous prediction accuracy to complex models across all prediction variables, and do so with substantially lower computational requirements. Table 2 provides the training time and inference time (time to generate forecasts) of the baseline models, and shows that simple models require roughly 500x less computation for inference. Hence, in practical systems the use of simple models would enable shorter prediction intervals, leading to higher frequency control, and allow for the use of lower cost compute hardware. The similar prediction accuracies of the tested models are found to lead to similar operational performance when used in the MPC scheme, see Fig. 2d, where the dashed line indicates the bound on operational performance achieved by an MPC scheme with perfect forecasts. The use of the MPC controllers with the specified battery systems leads to an average 8.7% improvement in operational performance for the multi-building energy system, with a range of 7.3% to 9.8%.
Model
Linear
Conv
ResMLP
TFT
NHiTS
DeepAR
Training time (hrs)
12.3
26.8
13.4
8.9
15.2
9.3
Prediction inference time (s)
29
73
146
15,222
15,289
74,991
Table 2: Computation times for baseline models for training on 8 years of data and performing inference for 2 year duration simulations.
@@ 3.2 Model generalisation
The generalisation of prediction models between buildings is tested by using the baseline models trained on data from each of the 15 buildings to forecast electrical load for every other building. Fig. 4 shows the distribution of relative forecast accuracies achieved by each model over all buildings (DeepAR is excluded due to excessive computational costs), where the y-axis plots the nRMSE forecast accuracy of the tested model (potentially trained on a different building) normalised by the forecast accuracy achieved by the model trained on data from the target building. The Linear model provides similar generalisation performance to NHiTS, and is substantially better than all other models.
@@ 3.3 Data efficiency
Data usage for developing forecasting models incurs cost from both its acquisition and the computation required to exploit it. However, greater data usage can increase forecast accuracy, resulting in lower operational cost of the energy system. Therefore, any measures to reduce data volumes must be traded-off against the corresponding decrease in prediction accuracy. This section explores a series of aspects of data efficiency and their effect on forecast accuracy. Subsequently, section 3.4 investigates the impact of forecast accuracy on operational performance of the MPC.
@@ 3.3.1 Volume of training data
Using greater volumes of training data, longer durations of historic measurements of buildings, can improve model prediction accuracy by avoiding over-fitting, improving temporal generalisability, but comes at a roughly linear cost. The trade-off between the volume of training data used, the combined duration of the train and validate datasets, and forecast accuracy is investigated by training prediction models on subsets of the building energy dataset of varying durations from 8 years to 3 months. For the complex models, at least 2 years of training data was required due to the use of temporal covariates. Fig. 6 plots the proportional improvement in prediction accuracy of each model (negative improvement shows worsening model accuracy) compared to baseline when training using different data volumes. Across a broad range of model architectures, reducing training data volumes down to 2 years has a limited impact on prediction accuracy, in some cases improving prediction accuracy. Trends differ between prediction variables, however for the Linear model, in the majority of cases there is a small prediction accuracy penalty between 8 and 1 years of training data, and then a rapid worsening of forecasting accuracy below 1 year. This indicates that when making data collection decisions to support prediction model development for building electrical energy systems, at most 2 years of measurement data should be gathered.
@@
Once building data has been gathered, there remains a question as to whether all of the available data should be used for training, or whether some data should be excluded to improve model prediction. For instance, if data that is non-representative of the present building load dynamics can be excluded, prediction accuracy can be improved alongside data efficiency. Change-point analysis can be used to screen training data for this purpose, by detecting changes in building load dynamics and excluding data preceeding the change from training. A change-point analysis using the BEAST algorithm zhao2019DetectingChangepointTrend was performed on the building load dataset, described in E. Points where changes in load trend were detected were used to screen the training data for 7 selected buildings. In order to allow a better analysis of the detected change-points, the training dataset was increased to 7 years (2010-2016), leaving 1 year of validation data (2017). Linear prediction models were trained using sections of the training data starting at each of the detected change-points and running to the end of 2016.
Fig. 6 shows the relative improvements in prediction accuracy of models trained on data screened using change-points, compared to the case where the full 7 years of available data is used, where the x-axis position indicates the timing of the change-point. For most buildings, screening the training data using change-points improves prediction accuracy, indicating that non-representative data is being removed from the training dataset. This suggests that ex-post training data screening can be used to improve model accuracy whilst reducing data usage and so computational cost. These results correspond with the findings of choi2023PerformanceEvaluationDeep , which shows that additional training data only improves prediction accuracy if it is sufficiently similar to the test data, and can reduce accuracy when insufficiently similar. However, the results of this study demonstrate an additional consideration, which is that prediction accuracy worsens substantially when insufficient training data volume is used. This suggests a trade-off between having enough data to avoid over-fitting, and removing non-representative data which causes over-generalisation of the model. Figures 6 & 6 indicate that at least 1 year of training data is required to achieve good prediction accuracy, after which only sufficiently similar data should be included in the training dataset. The reason for the difference in results compared to choi2023PerformanceEvaluationDeep is likely due to this study testing model prediction accuracy on 2 years of data, compared to 0.1 years in choi2023PerformanceEvaluationDeep . As a result, models in this study must capture the seasonality of building load dynamics, hence requiring at least 1 year of training data to achieve good prediction accuracy.
@@ 3.3.2 Data features
The variables that can be monitored in building energy systems are often well correlated, for example ambient temperature and electrical load. As a result, covariates can be used by prediction models to attempt to identify underlying links between the variables and improve forecast accuracy. However, the use of additional data variables (features) incurs a cost from both the collection and exploitation of the extra data, for instance the purchase of proprietary weather forecasts. The impact of feature selection, and the corresponding data usage, on forecast accuracy is investigated by comparing the prediction accuracy of models trained using varying numbers of data features. The Pearson correlation coefficient between the 11 data features available in the building energy dataset are shown for the case of Building 0 in Fig. 7. Linear prediction models were trained using the $n$ most correlated variables/features for each of the electrical load, solar generation, electricity price, and carbon intensity target variables. Fig. 8 shows the impact of data feature usage on forecast accuracy. For most prediction variables, the inclusion of additional data features in the model worsens forecast accuracy. Several factors could contribute to this behaviour:
@@
    Over-fitting hawkins2004ProblemOverfitting : The incorporation of an excessive number of features may lead the model to over-fit the training data, worsening its predictions for unseen data.
@@
    Multi-collinearity farrar1967MulticollinearityRegressionAnalysis : The presence of highly correlated features can destabilize the model, resulting in worse prediction accuracy.
@@
    Curse of Dimensionality verleysen2005CurseDimensionalityData : An increase in the number of features increases the dimensionality of the data, which can lead to data sparsity and degradation of prediction accuracy.
These results show that testing should be performed before decisions are made regarding the collection of covariate data to ensure that, for the type of system in question, its use will firstly improve prediction accuracy, and secondly that said accuracy improvements warrant the cost of the data.

### v2

@@ 2.2 Forecasting task & performance evaluation
For the MPC scheme used, Eq. 1, forecasts of the following operational condition variables over the planning horizon $T$ are required at each time instance, $t$:
@@
This study investigates the impacts of data on the forecasting of these 4 types of operational variables. The accuracies of these forecasts are quantified using two error metrics, the normalised Mean Absolute Error (nMAE), and the normalised Root Mean Squared Error (nRMSE), given by Eqns 2 & 3,
[…]
$\frac{\frac{1}{N}{\sum_{t=0}^{N-1}}\sqrt{\frac{1}{T}\sum_{\tau=1}^{T}\left(f^{v}_{t,\tau}-v_{t+\tau}\right)^{2}}}{\frac{1}{N}\sum_{t=0}^{N-1}v_{t}}$ (3)
@@
The operational performance achieved by the MPC scheme using a given set of prediction models is quantified by evaluating the objective specified in Eq. 1 with the simulated behaviour of the building energy system resulting from the use of the controller.
@@ 2.3 Cambridge Estates electricity use dataset
A dataset of historic building electricity usage measurements from a set of buildings across the Cambridge University Estates covering the period 2010 to 2019 [46] is used to provide a case study for the multi-building energy system. The dataset consists of 10 years of historic electricity usage data from 30 buildings of varying use types in the University Estate, such as lecture blocks, offices, laboratories, and museums, alongside weather observations and grid electricity price and carbon intensity data. The electricity usage measurements record the total electrical load of each building, which includes lighting, plug loads, and plant equipment electricity consumption. It is assumed that none of the buildings have heat pumps or AC units installed, meaning the electricity usage does not include any contributions from space heating or cooling.
In all, the dataset contains the following variables: building electrical load data, weather data for Cambridge from the Met Office MIDAS dataset [47] (temperature and relative humidity) and renewables.ninja reanalysis model [48, 49] (direct and diffuse solar irradiance), dynamic electricity pricing tariff data from Energy Stats UK [50], grid electricity carbon intensity from the National Grid ESO Data Portal [51], and temporal information including hour, day, and month indices, as well as daylight savings status. All data is available at hourly resolution. Further detail on the data sourcing and processing is available in reference [46].
@@ 2.4.1 Simple neural models
Recent literature, [30], has shown that simple neural models using Direct Multi-Step forecasting (DMS), where all predictions over the forecast window are generated concurrently, can outperform complex, transformer-based models using traditional, Iterated Multi-Step forecasting (IMS), in which a single-step forecaster is applied iteratively to generate a multi-step forecast. For all three simple neural architectures investigated, DMS forecasting is used.
@@ 2.4.2 State-of-the-art machine learning models
The Temporal Fusion Transformer (TFT) model [52], developed by Google, is an attention-based architecture that enables the fusion of data from multiple input sources to inform predictions. The neural structure contains features which allow for the learning of multiple underlying relationships across temporal scales, and the attention mechanism allows for interpretation of the model predictions, i.e. which data the model is exploiting to produce its forecasts. The model uses categorical covariates of date-time information, as well as temperature information for predicting building loads.
@@
Neural Hierarchical Interpolation for Time Series Forecasting (NHiTS) [53] is an MLP model which learns a set of basis functions at different frequencies that describe the underlying patterns in the training data, and produces forecasts by using hierarchical interpolation to combine predictions from the basis functions in a computationally efficient manner. It uses categorical covariates of date-time information.
@@
DeepAR is a Recurrent Neural Network (RNN) based model developed by Amazon [54], which has been widely applied in a range of research areas. It is a probabilistic forecasting model, but for this study only the mean prediction is used. The model uses categorical covariates of date-time information.
@@ 3.1 Baseline prediction accuracy comparison
The prediction accuracy of each forecasting model trained using 8 years of training data, the maximum available, was evaluated to provide a baseline for the model architectures in a setting without data limitations. For brevity, forecast accuracy results are discussed for the nRMSE metric only, however equivalent results were found with the nMAE metric. Figures 2a & 2b show that the simple Linear model achieves similar or better prediction accuracy (lower nRMSE values) compared to the complex, state-of-the-art models across all prediction variables. The Conv and ResMLP models provide similar accuracy when predicting building electrical loads, however both have significantly worse accuracy for the electricity price and carbon intensity prediction variables. Complex models achieve slightly better accuracy for electricity price and solar generation predictions, at most 8.5% and 4.2% lower nRMSE than the Linear model, achieved by NHiTS and TFT respectively. However, for some buildings the complex models exhibit very poor forecast accuracy for load predictions. These instances of poor accuracy are found to correlate with a measure of the similarity between the train and test datasets for building load, called the ‘Wasserstein similarity metric’, which is described in D and is used for the study of model reuse in the following section. Fig. 2c plots the relationship between prediction accuracy and data similarity for each model, and shows that complex models provide poor prediction accuracy when the training and test data are significantly different. Hence, simple neural models provide better prediction generalisation under changes in building load dynamics between the training and test data, which is highly advantageous for application to practical systems, as occupant driven load dynamics may change after system installation, e.g. due to a change of building use.
@@
These results indicate that simple neural models have sufficient expressivity to well capture the underlying trends in building energy usage, but learn sufficiently simple relationships about the data as to avoid the problem of over-fitting experienced by the complex models, leading to better generalisation in time. This suggests that the majority of the trends in the load data are present within the 1 week input window used by the simple models, which agrees with the strong daily and weekly trends of typical building energy usage, caused by building occupancy patterns. It is likely that annual trends are well captured by the mean value over the input window, as these trends are slow relative to the prediction length. Additionally, the predictions required by MPC are relatively short in length, 48 hours in this case, whereas the complex models are found to provide more stable predictions and greater overall accuracy for longer duration forecasts [53, 55].
Overall, simple neural models are able to provide analogous prediction accuracy to complex models across all prediction variables, and do so with substantially lower computational requirements. Table 2 provides the training time and inference time (time to generate forecasts) of the baseline models, and shows that simple models require roughly 500x less computation for inference. Hence, in practical systems the use of simple models would enable shorter prediction intervals, leading to higher frequency control, and allow for the use of lower cost compute hardware. Further, due to their high computational cost, the complex models were only able to use a 72 hour input window [56]. As a result, they have less information available to inform their predictions, likely limiting the accuracy they could achieve. Therefore, the computational efficiency of simple neural models enabling the use of longer input windows provides an additional advantage.
Model Linear Conv ResMLP TFT NHiTS DeepAR
Training time (hrs) 12.3 26.8 13.4 8.9 15.2 9.3
Prediction inference time (s) 29 73 146 15,222 15,289 74,991
Table 2: Computation times for baseline models, trained on 8 years of data, and predicting for simulations of 2 years duration.
The similar prediction accuracies of the tested models are found to lead to similar operational performance when used in the MPC scheme, see Fig. 2d, where the dashed line indicates the bound on operational performance achieved by an MPC scheme with perfect forecasts. The use of the MPC controllers with the specified battery systems leads to an average 8.7% improvement in operational performance for the multi-building energy system, with a range of 7.3% to 9.8%.
@@ 3.2 Model generalisation
The generalisation of prediction models between buildings is tested by using the baseline models trained on data from each of the 15 buildings to forecast electrical load for every other building. Fig. 4 shows the distribution of relative forecast accuracies achieved by each model over all buildings (DeepAR is excluded due to excessive computational costs), where the y-axis plots the nRMSE forecast accuracy of the tested model (potentially trained on a different building) normalised by the forecast accuracy achieved by the model trained on data from the target building. The Linear model provides similar generalisation performance to NHiTS, and is substantially better than all other models. It is proposed that the Linear model, being mechanistically equivalent to linear regression, achieves good generalisation as it learns relatively simple relationships about the load data, which are consistent between buildings. Whereas NHiTS achieves good generalisation due to its behaviour of generating smooth forecasts [53], making it less susceptible to producing erratic and inaccurate predictions with unseen input data. As before, TFT suffers from over-fitting, leading to relatively poor generalisation. Whilst the Conv model was able to provide good generalisation in time where the dataset similarity was close, indicated by small Wasserstein metric values in Fig. 2c, between buildings the load data is much less similar, and the relationships learned by the Conv model are no longer valid, leading to poor prediction accuracy. It is suggested this is due to the features learned by the pooling process no longer being pertinent for the new building.
@@ 3.3 Data efficiency
Data required for developing forecasting models incurs cost from both its acquisition and the computation required to exploit it. However, using longer durations of training data or additional data variables can increase forecast accuracy, resulting in lower operational cost of the energy system. Therefore, any measures to reduce the quantity of data used must be traded-off against the corresponding decrease in prediction accuracy. This section explores a series of aspects of data efficiency and their effect on forecast accuracy. Subsequently, section 3.4 investigates the impact of forecast accuracy on operational performance of the MPC.
@@ 3.3.1 Volume of training data
Using greater volumes of training data, longer durations of historic measurements of buildings, can improve model prediction accuracy by avoiding over-fitting, improving temporal generalisability, but comes at a roughly linear cost. The trade-off between the length of training data used, the combined duration of the train and validate datasets, and forecast accuracy is investigated by training prediction models on subsets of the building energy dataset of varying durations from 8 years, the maximum duration available, to 3 months (one season), the shortest duration tested in [28]. For the complex models, at least 2 years of training data was required due to the use of temporal covariates. Fig. 6 plots the proportional improvement in prediction accuracy of each model (negative improvement shows worsening model accuracy) compared to baseline when training using different data volumes. Across a broad range of model architectures, reducing training data length down to 2 years has a limited impact on prediction accuracy, in some cases improving prediction accuracy. Trends differ between prediction variables, however for the Linear model, in the majority of cases there is a small prediction accuracy penalty between 8 and 1 years of training data, and then a rapid worsening of forecasting accuracy below 6 months. This indicates that when making data collection decisions to support prediction model development for building electrical energy systems, at most 2 years of measurement data should be gathered. However, at least two seasons of data (6 months) are required to prevent model over-fitting and learn trends which generalise across seasons.
@@
Once building data has been gathered, there remains a question as to whether all of the available data should be used for training, or whether some data should be excluded to improve model prediction. For instance, if data that is non-representative of the present building load dynamics can be excluded, prediction accuracy can be improved alongside data efficiency. Change-point analysis can be used to screen training data for this purpose, by detecting changes in building load dynamics and excluding data preceding the change from training. A change-point analysis using the BEAST algorithm [57] was performed on the building load dataset, described in E. Points where changes in load trend were detected were used to screen the training data for 7 selected buildings. In order to allow a better analysis of the detected change-points, the training dataset was increased to 7 years (2010-2016), leaving 1 year of validation data (2017). Linear prediction models were trained using sections of the training data starting at each of the detected change-points and running to the end of 2016.
Fig. 6 shows the relative improvements in prediction accuracy of models trained on data screened using change-points, compared to the case where the full 7 years of available data is used, where the x-axis position indicates the timing of the change-point. For most buildings, screening the training data using change-points improves prediction accuracy, indicating that non-representative data is being removed from the training dataset. This suggests that ex-post training data screening can be used to improve model accuracy whilst reducing the quantity of training data used, and so computational cost. These results correspond with the findings of [28], which shows that additional training data only improves prediction accuracy if it is sufficiently similar to the test data, and can reduce accuracy when insufficiently similar. However, the results of this study demonstrate an additional consideration, which is that prediction accuracy worsens substantially when insufficient training data volume is used. This suggests a trade-off between having enough data to avoid over-fitting, and removing non-representative data which causes over-generalisation of the model. Figures 6 & 6 indicate that at least 1 year of training data is required to achieve good prediction accuracy, after which only sufficiently similar data should be included in the training dataset. The reason for the difference in results compared to [28] is likely due to this study testing model prediction accuracy on 2 years of data, compared to 0.1 years in [28]. As a result, models in this study must capture the seasonality of building load dynamics, hence requiring at least 1 year of training data to achieve good prediction accuracy.
@@
Change-point analysis can also be used online to detect real time changes in building load dynamics and provide indication of when the prediction models should be updated (i.e. retrained using recent data), due to the training data of the current model no longer being representative of the building behaviour, causing reduced prediction accuracy. This is particularly pertinent in the context of climate change, which is expected to significantly impact the energy usage behaviour of buildings, for instance reducing peak heating loads due to higher winter temperatures, and increasing summer electrical loads to provide cooling during prolonged heat waves.
@@ 3.3.2 Data features
The variables that can be monitored in building energy systems are often well correlated, for example ambient temperature and electrical load. As a result, covariates can be used by prediction models to attempt to identify underlying links between the variables and improve forecast accuracy. However, the use of additional data variables (features) incurs a cost from both the collection and exploitation of the extra data, for instance the purchase of proprietary weather forecasts. The impact of feature selection on forecast accuracy is investigated by comparing the prediction accuracy of models trained using varying numbers of data features. The Pearson correlation coefficient between the 11 data features available in the building energy dataset are shown for the case of Building 0 in Fig. 7. Linear prediction models were trained using the $n$ most correlated variables/features for each of the electrical load, solar generation, electricity price, and carbon intensity target variables. Feature selection was performed separately for each case by ranking Pearson correlation coefficients between the target variable and all other available covariates. Fig. 8 shows the impact of the number of data features used by the models on forecast accuracy. For most prediction variables, the inclusion of additional data features worsens forecast accuracy. Several factors could contribute to this behaviour:
@@
    Over-fitting [58] : The incorporation of an excessive number of features may lead the model to over-fit the training data, worsening its predictions for unseen data.
@@
    Multi-collinearity [59] : The presence of highly correlated features can destabilize the model, resulting in worse prediction accuracy.
@@
    Curse of Dimensionality [60] : An increase in the number of features increases the dimensionality of the data, which can lead to data sparsity and degradation of prediction accuracy.
These results show that testing should be performed before decisions are made regarding the collection of covariate data to ensure that, for the type of system in question, its use will firstly improve prediction accuracy, and secondly that said accuracy improvements warrant the cost of the data. This study considers a simple, correlation based feature selection method, however more advanced, optimization based techniques [61, 62] could be used to determine the set of available features which provides the optimal trade-off between data cost and model prediction accuracy.